OpenStudy (anonymous):

find the six trig values if given cos theta = 1/2 and sin theta < 0 need to check the sin theta I get negative square root of 3/4

OpenStudy (jtvatsim):

yes, but you can simplfy your answer for sin a little bit more...

OpenStudy (anonymous):

how?

OpenStudy (jtvatsim):

\[-\sqrt{\frac{3}{4}}=-\frac{\sqrt{3}}{\sqrt{4}}=-\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

thanks!